OpenStudy (anonymous):

Amy paid $94.12 for a pair of running shoes during a 15% off sale. What was the regular price. The regular price was_____ (Round off to the nearest cent, if necessary)

OpenStudy (anonymous):

110.7$

OpenStudy (anonymous):

Let x = unknown regular price x-0.15x=94.12 x(1-0.15)=94.12 0.85x=94.12 x=94.12/0.85 x=$110.73

OpenStudy (anonymous):

@Toyjia Do you need more help?

OpenStudy (anonymous):

Frances left a 21% tip on a meal.Including the tip, its cost was $39.93. What was the cost of the meal before the tip? The cost of the meal before the tip was $____ .

OpenStudy (anonymous):

0.21x+x=39.93 x(0.21+1)=39.93 x(1.21)=39.93 x=39.93/1.21 x=$33.00

OpenStudy (anonymous):

The estimated insurance loss from a hurricane was $24.7 billion. this was $14.7 billion more than the loss from a tsunami three years ago later. What was the estimated insurance loss from the tsunami? $____ billion Simplify your answer. type an integer or a decimal.